OpenStudy (anonymous):

-(85p + 67q + 40) = ??

OpenStudy (anonymous):

Come on guys....Please help me.

OpenStudy (moongazer):

-85p-67q-40

OpenStudy (anonymous):

85p+67q=-40

OpenStudy (moongazer):

there is no =0 and distribute first the negative sign

OpenStudy (anonymous):

OMG, thank you!

OpenStudy (anonymous):

i'm so sorry. i'm reding is wrong

OpenStudy (moongazer):

nuryamin is wrong i think it is -85p-67q-40=0 -85-67q=40 because you need to distribute the negative sign